# The empirical median for estimating the common mean of heteroscedastic random variables

### v1

@@ Abstract
We study the problem of mean estimation in the heteroscedastic setting. In particular, we consider symmetric random variables having the same location parameter and different and unknown scale parameters. Our goal is then to estimate their unknown common location parameter. It is an elementary topic but yet a not very well-studied one since we always make the assumption that the random variables are independent and identically distributed. In this paper, we study the median estimator and we establish upper and lower bounds on its estimation error that are of the same order and that generalize and improve recent results of Devroye et al. [7] and Xia [22].
@@ 1 Introduction
Let $(X_{1},\dots,X_{n})$ be $n$ real symmetric random variables having the same location parameter $\theta$ and different and unknown scale parameters $\sigma_{i}$, $i\in[\![1,n]\!]$. We make the assumption that the $\sigma_{i}$’s are ordered without loss of generality meaning that $0<\sigma_{1}\leqslant\dots\leqslant\sigma_{n}<+\infty$. However, the estimators we consider lack awareness of this order, and thus we assume that they are invariant to the permutation of the variances. We aim at constructing an estimator of $\theta$ taking advantage of the heteroscedastic structure of the problem. This problem can be considered as "dual" to robustness setting. Indeed, in robust estimation, one attempts to limit the damaging effect of a small number of outliers. In heteroscedastic estimation, one would like to exploit the possible benefit of a small number of highly informative observations. If the densities of the $X_{i}$’s and thus the scale parameters $\sigma_{i}$’s are known, the problem becomes easily tractable. The optimal estimator in this case is the maximum likelihood estimator, which is asymptotically minimax under suitable assumptions. However, the scale parameters are unknown and cannot be estimated, as only one observation of $\sigma_{i}$, is available for each $i\in[\![1,n]\!]$. The bound of the maximum likelihood estimator, which will be detailed later for the Gaussian case, is therefore overly optimistic and cannot be attained.
Notations
Throughout the paper, we use the following notation. We denote $M:=\mathrm{Med}(X)$ the theoretical median of a random variable $X$ and $\widehat{\mathrm{med}}(X_{1},\dots,X_{n})$ the empirical median. Moreover, $\lfloor$.$\rfloor$ . ⌋ denotes the floor function.
A naive way to tackle this problem in the Gaussian framework when $X_{i}\sim\mathcal{N}(\theta,\sigma_{i}^{2}),$ $i=1,\dots,n,$ is to study the empirical mean estimator which is defined by
@@
and which follows a Gaussian distribution with mean $\theta$ and variance $\sigma^{2}=\frac{1}{n^{2}}\sum_{i=1}^{n}\sigma^{2}_{i}$. And so, for $\delta\in(0,1)$, we have that with probability at least equal to $1-\delta$,
@@
It can be observed that this estimator is highly sensitive to observations with high variances, yet relatively insensitive to ones with small variances that contain most information about the mean. In fact, if we fix the first $n-1$ variances and let $\sigma_{n}$ tend to infinity, Bound (1) increases in this case. Conversely, if we assume that the first standard deviation $\sigma_{1}$ tends to zero, thereby implying a high degree of certainty regarding the observation in question, we observe that Bound (1) remains relatively stable.
@@
and which also follows a Gaussian distribution with mean $\theta$ and variance $\sigma^{2}=\Big{(}\sum_{i=1}^{n}\sigma_{i}^{-2}\Big{)}^{-1}$. So, for $\delta\in(0,1)$, we have with probability at least equal to $1-\delta$ that
@@
These two results are based a classical tail bound that can be found in [15, p. 9] which gives that if $X\sim\mathcal{N}(0,1)$, then for all $t>0$, $\mathbb{P}\left(\lvert X\rvert\geqslant t\right)\leqslant\exp\left(-\frac{t^{2}}{2}\right).$
Consequently, this estimator is highly responsive to minor fluctuations. Indeed, if the smallest standard deviation $\sigma_{1}$ tends to zero, Bound (2) also tends to zero. Furthermore, the estimator is not sensitive to observations with large variances, as demonstrated by the fact that when $\sigma_{n}$ tends to +$\infty$+ ∞, Bound (2) remains bounded. Moreover, and in the Gaussian case, the maximum likelihood estimator is minimax in the sense of the squared error [11, Section 16.4 p. 330]. However, this estimator is based on the assumption of prior knowledge of the variances, which is not applicable in our case.
Our main goal is therefore to construct an estimator that will be obviously better than the empirical mean and that will try to be as good as that of the maximum likelihood estimator provided by the oracle, bearing in mind that the latter can never be attained since the variances are unknown.
@@
The paper is organized as follows. We will first analyze the empirical median estimator in Section 2 by establishing and upper bound and a lower one that matches the upper bound to within a numerical constant. In Section 3, we first expose prior works and recent results and then show that our result on the upper bound improves and generalizes two of the previous works. And finally, we give in Section 4 the proofs of the established results.
@@ 2 Analysis of the empirical median estimator: Main results
Given the symmetry of the problem, the empirical median estimator is a natural one to study. Moreover, in contrast to the mean, the median is relatively unaffected by high variances. In this section, the objective is to control the probability of deviation of the median and to obtain upper and lower bounds on the empirical median in high probability.
@@ 2.1 Upper bound
We state the theorem that determines the upper bound of the error in estimating the empirical median.
@@
Let $\delta\in(0,1)$, $t\geqslant 0$ and $C$ a positive constant such that $10<C<$ 1 and $\delta>\exp{\left(-2nC^{2}\right)}$. Let $(X_{1},\dots,X_{n})$ be $n$ independent random variables that are symmetric around their mean $\theta$ and such that for all $i\in[\![1,n]\!]$ there exists $\sigma_{i}$ such that if $0\leqslant t\leqslant\sigma_{i}$ then
@@
Note that the condition in Theorem 1 is satisfied if $\frac{X_{i}}{\sigma_{i}}$ has a bounded density on some interval around $\theta$ and does not require the assumption of symmetry. This means that the condition of symmetric random variables around $\theta$ such that $\mathbb{P}(\theta\leqslant X_{i}\leqslant\theta+t)\geqslant C\frac{t}{\sigma_{i}}$ can be replaced by the following condition
@@
Remark
This corollary follows from Theorem 1 with $C=\frac{\exp{(-\frac{1}{2})}}{\sqrt{2\pi}}=0.24$.
@@ 2.2 Lower bound
As discussed in Section 1, no estimator can have an estimated error that is smaller than that of the maximum likelihood estimator, which is aware of all the variances and whose bound is minimax. However, in the absence of knowledge of these variances, the problem becomes significantly more challenging. It would therefore be of interest to identify a lower bound that is as close as possible to $\sqrt{\frac{2\log{(\frac{1}{\delta})}}{\sum_{i=1}^{n}\sigma_{i}^{-2}}}$.
We have the following lower bound on the estimation error of the empirical median estimator.
@@
The proof of Theorem 2 relies on Corollary 2 of the following lemma, the proof of which is given in Section 4.3.
@@
Let $\delta\in(\exp{(-n)},1)$ and $S=\sum_{i=1}^{n}V_{i}$, where $V_{1},\dots,V_{n}$ are independent random variables with $V_{i}\sim\mathcal{B}(p_{i})$ for all $i\in[\![1,n]\!]$. If $p_{i}\in[\frac{1}{4},\frac{3}{4}]$ for every $i\in[\![1,n]\!]$, then
@@
We conclude that the lower bound on the estimation error of the empirical median estimator matches the upper bound to within a numerical constant. Noteworthy, the smallest $\sqrt{n}$ values of $\sigma_{i}$ dot not influence the behavior of the median.
@@ 3 Discussion and prior work
Several foundational works laid the groundwork for robust estimation techniques. The seminal article by Huber [10] introduced robust estimators that mitigate the impact of outliers, a principle further extended by Tukey’s bi-weight estimator [21]. In the context of heteroscedasticity,  [2] analyzed weighted means and highlighted the limitations of traditional approaches when variances are only partially known. Building on median-based methods,  [17] proposed the Least Median of Squares estimator, which is particularly effective in settings with non-homogeneous variances and outliers where [3] introduced an estimator tailored for heavy-tailed distributions, leveraging advanced tail-bound techniques.
@@
A more recent study by [6] introduces a novel "balanced estimator." This estimator, a variation of modal estimators, imposes an additional condition: the number of observations in the right neighborhood of the estimator must approximately equal those in the left neighborhood. This balanced approach addresses asymmetries in data distributions and enhances estimator performance under specific conditions, showcasing the evolving landscape of robust mean estimation techniques.
@@
The upper bound on the estimation error of the empirical median that we established improves and generalizes on the results of Devroye, Lattanzi and Lugosi [7] and Xia [22]. Indeed, our bound is applicable in a more general setting than that of Xia [22] and is more accurate than that of Devroye et al. [7].
@@
while our estimator ensures this bound for any configuration of $\sigma_{1},\dots,\sigma_{n}$. Furthermore, our bound reveals the fact that the lowest $\sqrt{n}$ $\sigma_{i}$’s dot not influence the behavior of the median, while this fact is hidden when the condition (3) is imposed.
@@
which includes an additional logarithmic factor compared to our bound. Furthermore, the term under the maximum in (4) corresponding to $j=1$ yields the bound of [22]. But again, given that it is a maximum over $j$, our bound performs better overall. Moreover, the bound of [7] requires, in addition to the symmetry of the random variables, some tail assumptions on the distribution. In contrast, our bound is valid in a more general case requiring only the boundedness of the density around 00 of the random variables. Thus, our result is more accurate but also applicable in a more general setting than the one of [7].
@@
$\frac{\sqrt{\Big{(}\sum_{i=1}^{n}\sigma^{2}_{i}\Big{)}}}{n}\geqslant\frac{\sqrt{n}}{\sum_{i=j+1}^{n}\sigma_{i}^{-1}},$
where $j$ is a defined before namely, $j=\lfloor 2.93\sqrt{n\log{(\frac{1}{\delta})}}\rfloor$ which is equivalent to $\sqrt{n}$. We have that $\frac{\sqrt{\sum_{i=1}^{n}\sigma^{2}_{i}}}{\sqrt{n}}\geqslant\frac{n}{\sum_{i=1}^{n}\sigma_{i}^{-1}}$ and that $\frac{\sqrt{\sum_{\sqrt{n}\leqslant i\leqslant n}\sigma^{2}_{i}}}{\sqrt{n-\sqrt{n}}}\geqslant\frac{n-\sqrt{n}}{\sum_{\sqrt{n}\leqslant i\leqslant n}\sigma_{i}^{-1}}.$ So, it is sufficient to show that $\frac{\sqrt{\sum_{\sqrt{n}\leqslant i\leqslant n}\sigma^{2}_{i}}}{\sqrt{n-\sqrt{n}}}\leqslant\frac{\sqrt{\sum_{i=1}^{n}\sigma^{2}_{i}}}{\sqrt{n}}.$ For all $n\geqslant 4$, we have that $n-\sqrt{n}\geqslant\frac{n}{2}$. Moreover, we have that $\sum_{\sqrt{n}\leqslant i\leqslant n}\sigma^{2}_{i}\leqslant\sum_{i=1}^{n}\sigma^{2}_{i}$. By combining these two points, we obtain that
$\frac{\sum_{\sqrt{n}\leqslant i\leqslant n}\sigma^{2}_{i}}{n-\sqrt{n}}$ $\leqslant 2\frac{\sum_{\sqrt{n}\leqslant i\leqslant n}\sigma^{2}_{i}}{n}\leqslant 2\frac{\sum_{i=1}^{n}\sigma^{2}_{i}}{n}.$
We then have that
$\frac{n-\sqrt{n}}{\sum_{\sqrt{n}\leqslant i\leqslant n}\sigma_{i}^{-1}}$ $\leqslant\frac{\sqrt{\sum_{\sqrt{n}\leqslant i\leqslant n}\sigma^{2}_{i}}}{\sqrt{n-\sqrt{n}}}\leqslant\frac{\sqrt{2\sum_{i=1}^{n}\sigma^{2}_{i}}}{\sqrt{n}}.$
And since $\frac{n-\sqrt{n}}{\sum_{\sqrt{n}\leqslant i\leqslant n}\sigma_{i}^{-1}}\geqslant\frac{n}{2\sum_{\sqrt{n}\leqslant i\leqslant n}\sigma_{i}^{-1}},$ we finally obtain that $\frac{\sqrt{n}}{\sum_{\sqrt{n}\leqslant i\leqslant n}\sigma_{i}^{-1}}\leqslant\frac{2\sqrt{2}\sqrt{\sum_{i=1}^{n}\sigma^{2}_{i}}}{n}.$ Thus, we can conclude that the upper bound on the estimation error of the empirical median estimator is better than the one of the empirical mean estimator which is also its lower bound.
@@
In this paper, we obtained the upper bound that can be applied in a more general setting than that of Gaussian random variables. In fact, this is also the case for our lower bound result. Indeed, the only argument related to the Gaussian setting is that for all $i\in[\![1,n]\!]$, $\mathbb{P}\left(\theta\leqslant X_{i}\leqslant\theta+t\right)\leqslant\frac{t}{\sigma_{i}\sqrt{2\pi}}.$ Thus, this result on the lower bound extends to random variables with bounded densities around $\theta$. Thus, we can even consider random variables with infinite moments or also let the distribution have heavy tails, which was not covered by [7].
@@
The empirical median estimator is not optimal in a general framework as it is not sensitive to the $\sqrt{n}$ smallest variances, which contain most information about the unknown mean. Indeed, our lower bound on the estimation error does not depend on the $\sqrt{n}$ smallest variances. An interesting open problem is to find an estimator that would be sensitive to the smallest variances and thus achieves optimality among all estimators.

### v3

# Performance of the empirical median for location estimation in heteroscedastic settings

@@ Abstract
We investigate the performance of the empirical median for location estimation in heteroscedastic settings. Specifically, we consider independent symmetric real-valued random variables that share a common but unknown location parameter while having different and unknown scale parameters. Estimation under heteroscedasticity arises naturally in many practical situations and has recently attracted considerable attention. In this work, we analyze the empirical median as an estimator of the common location parameter and derive matching non-asymptotic upper and lower bounds on its estimation error. These results fully characterize the behavior of the empirical median in heteroscedastic settings, clarifying both its robustness and its intrinsic limitations and offering a precise understanding of its performance in modern settings where data quality may vary across sources.
@@ 1 Introduction
Let $(X_{1},\dots,X_{n})$ be independent real-valued random variables symmetric around a common location parameter $\theta$ but with potentially different and unknown scales $\sigma_{i}>0$, for $i\in\{1,\dots,n\}$. Without loss of generality, we suppose that the scales are ordered as $\sigma_{1}\leqslant\cdots\leqslant\sigma_{n}$, although we focus on estimators that are invariant to permutations of the indices and do not exploit this ordering. In this heteroscedastic setting, our goal is to estimate $\theta$ when only a single observation is available from each $X_{i}$, and no direct information on the $\sigma_{i}$’s is accessible. While several recent works, detailed in Section 3, have proposed estimators that outperform the empirical median [7], by leveraging assumptions such as local modality [10, 19] or carefully aggregating the most informative observations, this paper takes a different angle. Rather than competing with these methods, our aim is to sharpen the theoretical understanding of the empirical median as a natural benchmark procedure.
Two natural estimators that one may consider at first glance are the empirical mean and the maximum likelihood estimator (MLE). In particular, in the Gaussian case, one immediately obtains explicit results regarding their estimation errors. Indeed, in this framework, namely when $X_{i}\sim\mathcal{N}(\theta,\sigma_{i}^{2}),$ $i=1,\dots,n$, a straightforward, though potentially suboptimal, approach consists in using the empirical mean estimator, defined by
@@
and following a Gaussian distribution with mean $\theta$ and variance $\sigma^{2}=\frac{1}{n^{2}}\sum_{i=1}^{n}\sigma^{2}_{i}$. For $\delta\in(0,1)$, we have for this estimator that with probability at least equal to $1-\delta$,
@@
It can be observed that this estimator is highly sensitive to observations with high variances, yet relatively insensitive to ones with small variances that contain most information about the mean. In fact, if we fix the first n1n-1 variances and let $\sigma_{n}$ tend to infinity, Bound (1) increases in this case. Conversely, if we assume that the first standard deviation $\sigma_{1}$ tends to zero, thereby implying a high degree of certainty regarding the observation in question, we observe that Bound (1) remains relatively stable.
@@
and which also follows a Gaussian distribution with mean $\theta$ and variance $\sigma^{2}=\Big(\sum_{i=1}^{n}\sigma_{i}^{-2}\Big)^{-1}$. For $\delta\in(0,1)$, we have with probability at least equal to $1-\delta$ that
@@
These two previous results are based a classical tail bound that can be found in [18, p. 9] which gives that if $X\sim\mathcal{N}(0,1)$, then for all $t>0$, $\mathbb{P}\left(\lvert X\rvert\geqslant t\right)\leqslant\exp\left(-\frac{t^{2}}{2}\right).$
Consequently, this estimator is highly responsive to minor fluctuations. Indeed, if the smallest standard deviation $\sigma_{1}$ tends to zero, Bound (2) also tends to zero. Furthermore, the estimator is not sensitive to observations with large variances, as demonstrated by the fact that when $\sigma_{n}$ tends to $\infty$, Bound (2) remains bounded. Moreover, and in this Gaussian framework, the maximum likelihood estimator is minimax in the sense of the squared error [14, Section 16.4 p. 330]. However, this estimator is based on the assumption of prior knowledge of the variances, which is not applicable in our case.
Our main goal in this paper is to provide matching non-asymptotic upper and lower bounds on the deviation of the empirical median. This contributes to a precise performance characterization of a simple, natural, and widely-used estimator. The empirical median may indeed be suboptimal in various regimes, particularly when the signal is concentrated in the lowest-variance observations, but its robustness, simplicity, and parameter-free nature make it a relevant object of study in its own right.
Notations
Throughout the paper, we use the following notation. We denote $M:=\mathrm{Med}(X)$ the theoretical median of a random variable $X$ and $\widehat{\mathrm{med}}(X_{1},\dots,X_{n})$ the empirical median. Moreover, $\lfloor.\rfloor$ denotes the floor function.
@@
The paper is organized as follows. We will first analyze the empirical median estimator in Section 2 by establishing an upper bound and a lower one that matches the upper bound within a numerical constant. In Section 3, we first expose prior works and recent results and then show that our upper bound result improves and generalizes two of the previous works. Finally, in Section 4, we give the proofs of the established results.
@@ 2 Analysis of the empirical median estimator: Main results
Given the symmetry of the problem, the empirical median estimator is a natural one to study. Moreover, unlike the mean, the median is relatively unaffected by high variances. In this section, the objective is to control the probability of deviation of the median and to obtain upper and lower bounds on the empirical median in high probability.
@@ 2.1 Upper bound
We state the theorem that determines the upper bound of the estimation error of the empirical median.
@@
Let $\delta\in(0,1)$, $t\geqslant 0$ and $C$ a positive constant such that $0<C<1$ and $\delta>\exp{\left(-2nC^{2}\right)}$. Let $(X_{1},\dots,X_{n})$ be $n$ independent random variables that are symmetric around the location parameter $\theta$ and such that for all $i\in[\![1,n]\!]$ there exists $\sigma_{i}$ such that, if $0\leqslant t\leqslant\sigma_{i}$, then
@@
Note that the condition in Theorem 1 is satisfied if $\frac{X_{i}}{\sigma_{i}}$ has a bounded density on some interval around $\theta$ and does not require the assumption of symmetry. This means that the condition of symmetric random variables around $\theta$ such that $\mathbb{P}(\theta\leqslant X_{i}\leqslant\theta+t)\geqslant C\frac{t}{\sigma_{i}}$ can be replaced by the condition that $\theta$ is a median of $X$ and that
@@
Note that by the above condition, $\theta$ is in fact the unique median of $X$.
@@
This result follows from Theorem 1 with $C=\frac{\exp{(-\frac{1}{2})}}{\sqrt{2\pi}}=0.24$.
@@ 2.2 Lower bound
As discussed in Section 1, no estimator can have an estimated error that is smaller than that of the maximum likelihood estimator, which is aware of all the variances and whose bound is minimax. However, in the absence of knowledge of these variances, the problem becomes significantly more challenging. It would therefore be of interest to identify a lower bound that would be larger than $\sqrt{\frac{2\log{(\frac{1}{\delta})}}{\sum_{i=1}^{n}\sigma_{i}^{-2}}}$.
We establish the following lower bound on the estimation error of the empirical median estimator.
@@
The proof of Theorem 2 relies on Corollary 2 derived from the following lemma, the proof of which is given in Section 4.4.
@@
Let $\delta\in(\exp{(-n)},1)$ and $S=\sum_{i=1}^{n}V_{i}$, where $V_{1},\dots,V_{n}$ are independent random variables with $V_{i}\sim\mathcal{B}(p_{i})$ for all $i\in[\![1,n]\!]$. If $p_{i}\in(\frac{1}{4},\frac{3}{4})$ for every $i\in[\![1,n]\!]$, then
@@
We conclude that the lower bound on the estimation error of the empirical median estimator matches the upper bound within a numerical constant. Noteworthy, the smallest $\sqrt{n}$ values of $\sigma_{i}$ do not influence the behavior of the median.
@@ 3 Discussion and prior work
Several foundational works laid the groundwork for robust estimation techniques. The seminal article by Huber [13] introduced robust estimators that mitigate the impact of outliers, a principle further extended by Tukey’s biweight estimator [26]. In the context of heteroscedasticity,  [2] analyzed weighted means and highlighted the limitations of traditional approaches when variances are only partially known. Building on median-based methods,  [21] proposed the Least Median of Squares estimator, which is particularly effective in settings with non-homogeneous variances and outliers where [3] introduced an estimator tailored for heavy-tailed distributions, leveraging advanced tail-bound techniques.
@@
A more recent study by [6] introduces a novel "balanced estimator". This estimator, a variation of modal estimators, imposes an additional condition: the number of observations in the right neighborhood of the estimator must approximately equal those in the left neighborhood. This balanced approach addresses asymmetries in data distributions and enhances estimator performance under specific conditions, showcasing the evolving landscape of robust mean estimation techniques.
Another line of work considers the estimation of a location parameter under symmetric noise with unknown distribution. This approach, introduced by Stone [23], consists in first estimating the noise density using kernel methods and then performing a Newton step to maximize the resulting log-likelihood. This yields an estimator closely related to the maximum likelihood estimator. Building on this idea, [11] provide finite-sample optimality results, and show that the estimation error is governed by the Fisher information of the unknown noise distribution. Their analysis highlights the central role of Fisher information in the symmetric location model, even in the absence of full parametric knowledge.
Finally, the question of whether classical minimax lower bounds can be matched by actual estimators in finite-sample regimes has recently been addressed by [7] in the context of location estimation. In their work, the authors investigate whether the standard Le Cam-type lower bounds, based on binary hypothesis testing, can be attained by estimators under finite-sample constraints. They show that for symmetric mixtures of log-concave densities, the minimax rate can indeed be achieved (up to logarithmic factors). However, they also prove that for the broader class of symmetric unimodal distributions, no estimator can universally attain the two-point testing lower bound.
@@
The upper bound on the estimation error of the empirical median that we established improves and generalizes on the results of Devroye, Lattanzi and Lugosi [8] and Xia [27]. Indeed, our bound is applicable in a more general setting than that of Xia [27] and is more accurate than the one of Devroye et al. [8].
@@
while our estimator ensures this bound for any configuration of $\sigma_{1},\dots,\sigma_{n}$. Furthermore, our bound reveals the fact that the lowest $\sqrt{n}$ $\sigma_{i}$’s do not influence the behavior of the median, while this fact is hidden when the condition (3) is imposed.
@@
which includes an additional logarithmic factor compared to our bound. Furthermore, the term under the maximum in (4) corresponding to $j=1$ yields the bound of [27]. But again, given that it is a maximum over $j$, our bound performs better overall. Moreover, the bound of [8] requires, in addition to the symmetry of the random variables, some tail assumptions on the distribution. In contrast, our bound is valid in a more general case requiring only the boundedness of the density around 0 of the random variables. Thus, our result is more accurate but also applicable in a more general setting than the one of [8].
@@
$\frac{\sqrt{\Big(\sum_{i=1}^{n}\sigma^{2}_{i}\Big)}}{n}\geqslant\frac{\sqrt{n}}{\sum_{i=j+1}^{n}\sigma_{i}^{-1}}.$
For that, we apply the following lemma.
Lemma 2.
Let $\sigma_{1}\leqslant\sigma_{2}\leqslant\dots\leqslant\sigma_{n}$ be positive real numbers. Then, for any integer $j\leqslant\lfloor n/2\rfloor-1$, we have
$\frac{\sqrt{\sum_{i=1}^{n}\sigma_{i}^{2}}}{n}\geqslant c\cdot\frac{\sqrt{n}}{\sum_{i=j+1}^{n}\sigma_{i}^{-1}},$
for some absolute constant $c>0$.
Proof.
Let $k:=\lfloor n/2\rfloor$. At least $n-k\geqslant n/2$ of the $\sigma_{i}$ satisfy $\sigma_{i}\geqslant\sigma_{k}$, then
$\sum_{i=1}^{n}\sigma_{i}^{2}\geqslant\sum_{i=k+1}^{n}\sigma_{i}^{2}\geqslant(n-k)\sigma_{k}^{2}\geqslant\frac{n}{2}\sigma_{k}^{2},$
which implies that
$\frac{\sqrt{\sum_{i=1}^{n}\sigma_{i}^{2}}}{n}\geqslant\frac{\sqrt{n\sigma_{k}^{2}/2}}{n}=\frac{\sigma_{k}}{\sqrt{2n}}.$
On the other hand, since the sequence is non-decreasing and $j+1\leqslant k$, we have that $\sigma_{j+1}\leqslant\sigma_{k}$ which implies that $\sigma_{j+1}^{-1}\geqslant\sigma_{k}^{-1}$. There are njn-j terms in the sum $\sum_{i=j+1}^{n}\sigma_{i}^{-1}$, and each is at most $\sigma_{j+1}^{-1}$, hence
$\sum_{i=j+1}^{n}\sigma_{i}^{-1}\leqslant(n-j)\,\sigma_{j+1}^{-1}\leqslant n\,\sigma_{j+1}^{-1}.$
Thus
$\frac{\sqrt{n}}{\sum_{i=j+1}^{n}\sigma_{i}^{-1}}\leqslant\frac{\sqrt{n}}{n\sigma_{j+1}^{-1}}=\frac{\sigma_{j+1}}{\sqrt{n}}\leqslant\frac{\sigma_{k}}{\sqrt{n}}.$
Combining both bounds, we obtain that
$\frac{\sqrt{\sum_{i=1}^{n}\sigma_{i}^{2}}}{n}\geqslant\frac{\sigma_{k}}{\sqrt{2n}}\geqslant\frac{1}{\sqrt{2}}\cdot\frac{\sigma_{k}}{\sqrt{n}}\geqslant\frac{1}{\sqrt{2}}\cdot\frac{\sqrt{n}}{\sum_{i=j+1}^{n}\sigma_{i}^{-1}}.$
Therefore, the claimed inequality holds with constant $c=1/\sqrt{2}$. ∎
@@
In this paper, we established an upper bound on the estimation error of the empirical median that is also valid in a more general setting than that of Gaussian random variables. In fact, this is also the case for our lower bound result. In fact, the only argument related to the Gaussian setting is that for all $i\in[\![1,n]\!]$, $\mathbb{P}\left(\theta\leqslant X_{i}\leqslant\theta+t\right)\leqslant\frac{t}{\sigma_{i}\sqrt{2\pi}}.$ Thus, this result on the lower bound extends to random variables with bounded densities around their median $\theta$. Thus, we can even consider random variables with infinite moments or also let the distribution have heavy tails, which was not covered by [8].
@@
The empirical median estimator is not optimal in a general heteroscedastic framework, as it fails to exploit the information carried by the $\sqrt{n}$ smallest variances. Our bounds reflect this limitation, since they remain unaffected by these variances. This shows that, although the empirical median is robust and simple, it cannot reach classical minimax optimality in such settings. Recently, this broader question has been clarified in the context of location estimation by [7], who investigated whether classical minimax lower bounds can be attained by finite-sample estimators. They showed that, while minimax rates can indeed be achieved (up to logarithmic factors) for symmetric mixtures of log-concave densities, no estimator can universally attain the two-point testing lower bound for the broader class of symmetric unimodal distributions. Altogether, our findings illustrate the fundamental limits of the empirical median and highlight open directions for developing estimators that can match minimax lower bounds in general frameworks.